The PH of 0.65 M Ca(OH)2, calcium hydroxide, is ? 0.11 0.19 13.81 13.89 12.00
Answer #1

Ca (OH)2 Cat² + 2014 [OH-] = 2x0-65 = 1-3M porna - loglon Et log [4:3] kw 2 3 [HT [on 1014 = [H+] 1. 3 13. CH+] = 7:69x10 bH